A charge q is to be divided into two small parts what should be the value of the charges on the parts so that force between the parts will be maximum

Grade:12

1 Answers

Vansh Chouhan
7 years ago
Let Q and (q-Q) be the 2 parts of total charge..
let r be the distance b/w the 2 parts..
F= 1/4pie epsilon x Q(q-Q)/ r^2
or F= k Q(q-Q)
where k= 1/4pie epsilon x r^2 = constant
now F= k (Qq- Q^2)
dF/dQ = k (q- 2Q)
for maxima dF/dQ = 0
so kq- 2kQ = 0
or kq= 2kQ
or Q = q/2
the 2 parts are q/2 and q/2